# Cyber security of OT networks: A tutorial and overview

### v1

@@ 3.2.2 Security considerations in OT networks
    Insecure Communication Protocols : Many industrial communication protocols used by PLCs, such as Modbus, DNP3, and EtherNet/IP, were not designed with security in mind. These protocols often lack encryption, authentication, or integrity checks, making them vulnerable to Eavesdropping ( Attackers can intercept unencrypted traffic), Man-in-the-Middle (MitM) Attacks (attackers can modify communication between the PLC and control systems) and replay Attacks (legitimate commands can be captured and replayed later by an attacker). There are protocols with built-in encryption and authentication (e.g., OPC UA, TLS), It needs to be adapted for making communication more robust and secure[8].
@@
    Firmware Vulnerabilities: Some PLCs allow firmware updates without verifying the authenticity of the firmware, enabling attackers to inject malicious code. Some manufacturers unintentionally include backdoors in their firmware, providing attackers with hidden access (example firmware occurred with Siemens’ S7-1200 and S7-1500 PLCs in 2013. Siemens included a special ”engineering mode” in the firmware, designed to allow factory engineers and service technicians to access the PLC in case of emergencies or maintenance needs. While this feature was convenient for Siemens personnel, it effectively functioned as a backdoor, allowing privileged access to the PLC, bypassing regular authentication measures. [9].
@@
    Stuxnet: Stuxnet was one of the first highly sophisticated worm that specifically targeted Siemens PLCs used in Iran’s nuclear facilities [10]. The worm infected Windows computers and then targeted Siemens Step7 software running on programmable logic controllers (PLCs) to reprogram it, causing the centrifuges to spin out of control while reporting normal operations to operators [11]. Stuxnet exploited four zero-day vulnerabilities in Microsoft Windows to spread and infect systems. Zero-day vulnerabilities are previously unknown and unpatched flaws in software. The worm spread via removable USB drives, exploiting the autorun feature to execute itself when the drive was connected to a computer. It also spread through network shares and used a variety of techniques to escalate privileges and gain administrative access on infected machines. Once inside a network, Stuxnet searched for Siemens Step7 software and PLCs and injected malicious code into the PLCs, which allowed it to manipulate the industrial processes they controlled. The worm targeted specific models of Siemens PLCs that controlled centrifuges used for uranium enrichment. The worm had sophisticated rootkit capabilities that allowed it to hide its presence on infected systems[12]. It intercepted and altered communication between the PLCs and the monitoring software, ensuring that operators remained unaware of the sabotage. It subtly altered the speeds of the centrifuges, causing them to spin at dangerous levels and ultimately leading to physical damage, while reporting normal operations to the monitoring systems. It caused significant damage ( approximately 1,000 centrifuges) to Iran’s nuclear program, showcasing the potential for cyber-physical sabotage [13]. Stuxnet demonstrated the potential for cyber weapons to achieve strategic military objectives and set a precedent for the development of Advanced Persistent Threats (APTs) [14] that target critical infrastructure.
@@
    Industroyer/CrashOverride (2016): Industroyer [15] targeted Ukraine’s power grid, causing widespread outages. The malware included modules to control industrial processes via ICS protocols, including IEC 104, IEC 61850, and OPC, to manipulate switches and circuit breakers [16]. These protocols are used in industrial control systems (ICS) for telecontrol and telemonitoring operate over TCP/IP networks. The malware uses a backdoor to maintain persistent access to the compromised network. This backdoor allows the attackers to issue commands to the infected systems remotely. To cover its tracks and hinder recovery efforts, Industroyer contains a data wiper module that erases crucial system files, rendering the infected systems inoperable. The initial infection vector for Industroyer is not fully known, but it is believed to have involved spear-phishing emails or other social engineering techniques to gain an initial foothold in the target network [17]. Industroyer’s payload modules are capable of communicating directly with ICS protocols, enabling the malware to send malicious commands to industrial equipment, such as opening circuit breakers to disrupt power flow [18]. The attack caused a power outage in parts of Kiev, affecting tens of thousands of residents for about an hour. It demonstrated the capability to disrupt critical infrastructure by directly interacting with PLCs and other ICS components.
@@
    Triton/Trisis (2017): Triton malware targeted Schneider Electric’s Triconex safety instrumented systems (SIS) used in critical infrastructure [19]. Triton was discovered in 2017 after it was used to attack a petrochemical plant in Saudi Arabia. Triton consists of several modules that interact with the Triconex SIS controllers [20]. These modules are capable of reading and writing to the controllers. It aimed to reprogram SIS controllers to either shut down operations or cause unsafe conditions. The malware installs a backdoor on the compromised system, allowing remote operators to execute commands and manipulate the SIS devices. Triton includes an execution framework that ensures the payload is delivered and executed on the target devices, and it can persist across reboots. The initial infection vector for Triton remains unclear, but it is believed to involve compromising the engineering workstation connected to the SIS network. Once inside the network, the attackers use various techniques to move laterally and gain access to the SIS controllers. It highlighted the risk of attacks on safety systems designed to prevent catastrophic failures in industrial environments.
@@ 4 Emerging Trends in OT Cybersecurity
As industrial environments continue to evolve, the emergence of new cybersecurity trends provides robust solutions to address vulnerabilities in OT systems. These trends leverage advanced technologies such as artificial intelligence (AI) [21], blockchain [22], and digital twins [23], improving the security posture of OT networks while addressing the unique challenges of IT-OT convergence. With OT systems playing critical roles in sectors such as manufacturing, energy, and transportation, the integration of these innovations is essential to protect against increasingly sophisticated cyber threats.
Unlike traditional IT networks, OT systems must balance real-time operational requirements, legacy infrastructure, and stringent uptime demands, which complicates their cybersecurity needs. Emerging technologies address these challenges by enabling real-time threat detection, improving incident response efficiency, and improving system resilience. For example, AI models trained on historical sensor data can predict anomalies in industrial processes [24], while blockchain ensures tamper-proof logging of system events [25]. Similarly, digital twins allow organizations to simulate cyberattack scenarios [26], providing insight into vulnerabilities without disrupting live operations.
Moreover, with the convergence of IT and OT systems through Industry 4.0 initiatives, cybersecurity measures must now account for an expanded attack surface [27]. Emerging trends focus on addressing vulnerabilities stemming from insecure legacy protocols, insufficient authentication mechanisms, and increased reliance on resource-constrained IoT devices. By adopting these advanced technologies, organizations can improve the reliability, safety, and resilience of their critical infrastructure, ensuring operational continuity in the face of evolving cyber threats. Below emerging trends demonstrate how innovative technologies are addressing the evolving challenges of OT cybersecurity, offering practical solutions for securing critical infrastructure. As these technologies mature, their adoption across industries will be pivotal in enhancing the safety and reliability of OT systems.
@@ 4.1 AI and ML for Threat Detection and automated Cyber defense
AI and ML are transforming industries by enabling automation, intelligence, and efficiency. In self-driving cars, AI and ML power real-time sensor fusion, object detection, and route optimization, ensuring safer and more efficient autonomous navigation. In underwater vehicle design [28][29] [30], they enhance autonomous operation by optimizing sensor placement, improving environmental adaptation, and enabling intelligent decision-making for deep-sea exploration. In healthcare, AI-driven Computer-Aided Diagnosis is revolutionizing medical imaging, detecting diseases like malaria with higher accuracy [31] and speed, assisting doctors in early diagnosis and treatment planning.
In the energy sector, it also holds boundless potential, enhancing power grid management through real-time monitoring and adaptive load balancing, improving renewable energy forecasting for better integration of solar and wind power, and enabling predictive maintenance in oil and gas to minimize downtime and operational risks. In recent years, it has also been utilized to strengthen cybersecurity in OT networks, analyzing vast real-time data from sensors, PLCs, and SCADA systems to enable anomaly detection and predictive threat modeling, helping prevent cyber threats before they escalate. Duke Energy, a major power utility in the US, has deployed ML models to analyze sensor data from its power grid [32]. These models detect deviations in parameters such as voltage and frequency, flagging potential cyberattacks or equipment failures before they occur. This approach has significantly improved the reliability and safety of its energy distribution systems. According to a 2023 article, Duke Energy’s hybrid model of engineered analytics and ML has proven to be an excellent but imperfect tool — more accurate than either pure AI/ML tools or engineered analytics alone [33]. AI is also being used in predictive maintenance [34], where abnormal vibration data from turbine sensors is analyzed to forecast equipment failures, minimizing downtime and reducing costs [35].
@@ 4.2 Zero Trust Architecture in OT Systems
Zero Trust Architecture (ZTA) principles, which enforce continuous verification of identity and access at every network segment, are now being adapted to OT environments [36]. This approach ensures that access to critical OT systems is restricted based on roles and credentials, significantly reducing the risk of unauthorized activities.
Siemens has integrated ZTA into its industrial automation products, implementing strict role-based access control (RBAC) [37] and multi-factor authentication in SCADA and PLC systems. For instance, Siemens’ SIMATIC controllers now include access management features that require continuous user verification and monitor actions performed by operators. This ensures that only authorized personnel can modify critical system parameters, protecting against insider threat.
@@ 4.3 Blockchain for Secure Event Logging
Blockchain technology is gaining traction in OT networks as a decentralized and tamper-proof mechanism for event logging [38]. By recording system logs in an immutable ledger, blockchain ensures the integrity of data and provides a robust audit trail for critical system changes. ABB, a global technology leader, employs blockchain-based solutions in its power distribution systems to log firmware updates and operational events [39]. This prevents tampering with logs or unauthorized firmware changes, ensuring compliance with cybersecurity standards. Additionally, the use of blockchain helps ABB’s systems maintain transparency and trust during audit.
@@ 4.4 Digital Twins for Proactive Security
Digital twin technology is being adopted to enhance cybersecurity in OT environments by creating real-time virtual replicas of physical systems [40]. These replicas enable organizations to simulate cyberattack scenarios and test defensive measures without disrupting live operations.
The Port of Rotterdam, Europe’s largest port, has implemented a digital twin to secure its complex OT network that oversees cargo logistics and terminal operations [41] [42]. The system simulates ransomware attacks on its SCADA systems, enabling the port’s security team to analyze vulnerabilities and refine incident response plans. This proactive approach helps ensure uninterrupted port operations while addressing potential threats.
@@ 4.5 Advanced Encryption for Resource-Constrained Devices
Securing data communication in resource-constrained OT devices, such as sensors and RTUs, requires lightweight encryption techniques that do not compromise device performance. Schneider Electric employs lightweight encryption algorithms in its EcoStruxure architecture, which is widely used in energy management systems [43]. These algorithms protect data transmitted between sensors and centralized control systems, ensuring the integrity of real-time operational data. The Transport Layer Security (TLS) 1.3 protocol, introduced in 2018, has been adopted for secure communication in power grids, safeguarding against data interception and tampering [44].
@@ 4.6 AI-Driven Incident Response
AI-driven incident response systems in OT environments leverage a combination of machine learning, anomaly detection, and automated decision-making algorithms to enhance cybersecurity resilience [45]. These systems employ advanced techniques like behavioral anomaly detection, where machine learning models continuously monitor device activity and network traffic for deviations from baseline patterns[46]. In addition to isolating compromised devices, they integrate with network segmentation protocols to minimize the spread of attacks. Sophisticated orchestration frameworks enable the automatic deployment of mitigation actions, such as blocking specific IP addresses or rerouting traffic, and rollback mechanisms use secure snapshots to restore devices to their pre-incident states [47]. In some cases, systems also use reinforcement learning to adapt and improve response strategies over time, making them more effective in handling novel, previously unseen attack vectors.
@@ 4.8 Integration of Cyber-Physical Security
Integrating cyber and physical security measures is essential for holistic OT protection [48]. Advanced systems now combine physical surveillance data, such as access logs, with cybersecurity monitoring to detect coordinated attacks.

### v2

@@ 3.2.2 Security considerations in OT networks
    Insecure Communication Protocols : Many industrial communication protocols used by PLCs, such as Modbus, DNP3, and EtherNet/IP, were not designed with security in mind. These protocols often lack encryption, authentication, or integrity checks, making them vulnerable to Eavesdropping ( Attackers can intercept unencrypted traffic), Man-in-the-Middle (MitM) Attacks (attackers can modify communication between the PLC and control systems) and replay Attacks (legitimate commands can be captured and replayed later by an attacker). There are protocols with built-in encryption and authentication (e.g., OPC UA, TLS), It needs to be adapted for making communication more robust and secure[s7threats].
@@
    Firmware Vulnerabilities: Some PLCs allow firmware updates without verifying the authenticity of the firmware, enabling attackers to inject malicious code. Some manufacturers unintentionally include backdoors in their firmware, providing attackers with hidden access (example firmware occurred with Siemens’ S7-1200 and S7-1500 PLCs in 2013. Siemens included a special ”engineering mode” in the firmware, designed to allow factory engineers and service technicians to access the PLC in case of emergencies or maintenance needs. While this feature was convenient for Siemens personnel, it effectively functioned as a backdoor, allowing privileged access to the PLC, bypassing regular authentication measures. [enlyze].
@@
    Stuxnet: Stuxnet was one of the first highly sophisticated worm that specifically targeted Siemens PLCs used in Iran’s nuclear facilities [Langner]. The worm infected Windows computers and then targeted Siemens Step7 software running on programmable logic controllers (PLCs) to reprogram it, causing the centrifuges to spin out of control while reporting normal operations to operators [kushner2013real]. Stuxnet exploited four zero-day vulnerabilities in Microsoft Windows to spread and infect systems. Zero-day vulnerabilities are previously unknown and unpatched flaws in software. The worm spread via removable USB drives, exploiting the autorun feature to execute itself when the drive was connected to a computer. It also spread through network shares and used a variety of techniques to escalate privileges and gain administrative access on infected machines. Once inside a network, Stuxnet searched for Siemens Step7 software and PLCs and injected malicious code into the PLCs, which allowed it to manipulate the industrial processes they controlled. The worm targeted specific models of Siemens PLCs that controlled centrifuges used for uranium enrichment. The worm had sophisticated rootkit capabilities that allowed it to hide its presence on infected systems[farwell2011stuxnet]. It intercepted and altered communication between the PLCs and the monitoring software, ensuring that operators remained unaware of the sabotage. It subtly altered the speeds of the centrifuges, causing them to spin at dangerous levels and ultimately leading to physical damage, while reporting normal operations to the monitoring systems. It caused significant damage ( approximately 1,000 centrifuges) to Iran’s nuclear program, showcasing the potential for cyber-physical sabotage [CCDCOE]. Stuxnet demonstrated the potential for cyber weapons to achieve strategic military objectives and set a precedent for the development of Advanced Persistent Threats (APTs) [stojanovic2020apt] that target critical infrastructure.
@@
    Industroyer/CrashOverride (2016): Industroyer [kozak2023industroyer] targeted Ukraine’s power grid, causing widespread outages. The malware included modules to control industrial processes via ICS protocols, including IEC 104, IEC 61850, and OPC, to manipulate switches and circuit breakers [ESET]. These protocols are used in industrial control systems (ICS) for telecontrol and telemonitoring operate over TCP/IP networks. The malware uses a backdoor to maintain persistent access to the compromised network. This backdoor allows the attackers to issue commands to the infected systems remotely. To cover its tracks and hinder recovery efforts, Industroyer contains a data wiper module that erases crucial system files, rendering the infected systems inoperable. The initial infection vector for Industroyer is not fully known, but it is believed to have involved spear-phishing emails or other social engineering techniques to gain an initial foothold in the target network [dragos]. Industroyer’s payload modules are capable of communicating directly with ICS protocols, enabling the malware to send malicious commands to industrial equipment, such as opening circuit breakers to disrupt power flow [sans]. The attack caused a power outage in parts of Kiev, affecting tens of thousands of residents for about an hour. It demonstrated the capability to disrupt critical infrastructure by directly interacting with PLCs and other ICS components.
@@
    Triton/Trisis (2017): Triton malware targeted Schneider Electric’s Triconex safety instrumented systems (SIS) used in critical infrastructure [mekdad2021threat]. Triton was discovered in 2017 after it was used to attack a petrochemical plant in Saudi Arabia. Triton consists of several modules that interact with the Triconex SIS controllers [trisis_drago]. These modules are capable of reading and writing to the controllers. It aimed to reprogram SIS controllers to either shut down operations or cause unsafe conditions. The malware installs a backdoor on the compromised system, allowing remote operators to execute commands and manipulate the SIS devices. Triton includes an execution framework that ensures the payload is delivered and executed on the target devices, and it can persist across reboots. The initial infection vector for Triton remains unclear, but it is believed to involve compromising the engineering workstation connected to the SIS network. Once inside the network, the attackers use various techniques to move laterally and gain access to the SIS controllers. It highlighted the risk of attacks on safety systems designed to prevent catastrophic failures in industrial environments.
@@ 4 Emerging Trends in OT Cybersecurity
As industrial environments continue to evolve, the emergence of new cybersecurity trends provides robust solutions to address vulnerabilities in OT systems. These trends leverage advanced technologies such as artificial intelligence (AI) [sarker2024ai], blockchain [gimenez2021achieving], and digital twins [holmes2021digital], improving the security posture of OT networks while addressing the unique challenges of IT-OT convergence. With OT systems playing critical roles in sectors such as manufacturing, energy, and transportation, the integration of these innovations is essential to protect against increasingly sophisticated cyber threats.
Unlike traditional IT networks, OT systems must balance real-time operational requirements, legacy infrastructure, and stringent uptime demands, which complicates their cybersecurity needs. Emerging technologies address these challenges by enabling real-time threat detection, improving incident response efficiency, and improving system resilience. For example, AI models trained on historical sensor data can predict anomalies in industrial processes [kashpruk2023time], while blockchain ensures tamper-proof logging of system events [bajramovic2019secure]. Similarly, digital twins allow organizations to simulate cyberattack scenarios [mustofa2024analyzing], providing insight into vulnerabilities without disrupting live operations.
Moreover, with the convergence of IT and OT systems through Industry 4.0 initiatives, cybersecurity measures must now account for an expanded attack surface [george2024impact]. Emerging trends focus on addressing vulnerabilities stemming from insecure legacy protocols, insufficient authentication mechanisms, and increased reliance on resource-constrained IoT devices. By adopting these advanced technologies, organizations can improve the reliability, safety, and resilience of their critical infrastructure, ensuring operational continuity in the face of evolving cyber threats. Below emerging trends demonstrate how innovative technologies are addressing the evolving challenges of OT cybersecurity, offering practical solutions for securing critical infrastructure. As these technologies mature, their adoption across industries will be pivotal in enhancing the safety and reliability of OT systems.
@@ 4.1 AI and ML for Threat Detection and automated Cyber defense
AI and ML are transforming industries by enabling automation, intelligence, and efficiency. In self-driving cars, AI and ML power real-time sensor fusion, object detection, and route optimization, ensuring safer and more efficient autonomous navigation. In underwater vehicle design [vardhan2024sample][vardhan2024fusion] [vardhan2022deep], they enhance autonomous operation by optimizing sensor placement, improving environmental adaptation, and enabling intelligent decision-making for deep-sea exploration. In healthcare, AI-driven Computer-Aided Diagnosis is revolutionizing medical imaging, detecting diseases like malaria with higher accuracy [kumar2023malaria] and speed, assisting doctors in early diagnosis and treatment planning.
In the energy sector, it also holds boundless potential, enhancing power grid management through real-time monitoring and adaptive load balancing, improving renewable energy forecasting for better integration of solar and wind power, and enabling predictive maintenance in oil and gas to minimize downtime and operational risks. In recent years, it has also been utilized to strengthen cybersecurity in OT networks, analyzing vast real-time data from sensors, PLCs, and SCADA systems to enable anomaly detection and predictive threat modeling, helping prevent cyber threats before they escalate. Duke Energy, a major power utility in the US, has deployed ML models to analyze sensor data from its power grid [akeenhancing]. These models detect deviations in parameters such as voltage and frequency, flagging potential cyberattacks or equipment failures before they occur. This approach has significantly improved the reliability and safety of its energy distribution systems. According to a 2023 article, Duke Energy’s hybrid model of engineered analytics and ML has proven to be an excellent but imperfect tool — more accurate than either pure AI/ML tools or engineered analytics alone [tdworld]. AI is also being used in predictive maintenance [jambol2024transforming], where abnormal vibration data from turbine sensors is analyzed to forecast equipment failures, minimizing downtime and reducing costs [ahadov2024predictive].
@@ 4.2 Zero Trust Architecture in OT Systems
Zero Trust Architecture (ZTA) principles, which enforce continuous verification of identity and access at every network segment, are now being adapted to OT environments [syed2022zero]. This approach ensures that access to critical OT systems is restricted based on roles and credentials, significantly reducing the risk of unauthorized activities.
Siemens has integrated ZTA into its industrial automation products, implementing strict role-based access control (RBAC) [russoindustrial] and multi-factor authentication in SCADA and PLC systems. For instance, Siemens’ SIMATIC controllers now include access management features that require continuous user verification and monitor actions performed by operators. This ensures that only authorized personnel can modify critical system parameters, protecting against insider threat.
@@ 4.3 Blockchain for Secure Event Logging
Blockchain technology is gaining traction in OT networks as a decentralized and tamper-proof mechanism for event logging [javed2023blockchain]. By recording system logs in an immutable ledger, blockchain ensures the integrity of data and provides a robust audit trail for critical system changes. ABB, a global technology leader, employs blockchain-based solutions in its power distribution systems to log firmware updates and operational events [baggio2020blockchain]. This prevents tampering with logs or unauthorized firmware changes, ensuring compliance with cybersecurity standards. Additionally, the use of blockchain helps ABB’s systems maintain transparency and trust during audit.
@@ 4.4 Digital Twins for Proactive Security
Digital twin technology is being adopted to enhance cybersecurity in OT environments by creating real-time virtual replicas of physical systems [wang2023survey]. These replicas enable organizations to simulate cyberattack scenarios and test defensive measures without disrupting live operations.
The Port of Rotterdam, Europe’s largest port, has implemented a digital twin to secure its complex OT network that oversees cargo logistics and terminal operations [wang2021multi] [brull2024role]. The system simulates ransomware attacks on its SCADA systems, enabling the port’s security team to analyze vulnerabilities and refine incident response plans. This proactive approach helps ensure uninterrupted port operations while addressing potential threats.
@@ 4.5 Advanced Encryption for Resource-Constrained Devices
Securing data communication in resource-constrained OT devices, such as sensors and RTUs, requires lightweight encryption techniques that do not compromise device performance. Schneider Electric employs lightweight encryption algorithms in its EcoStruxure architecture, which is widely used in energy management systems [abir2024industry]. These algorithms protect data transmitted between sensors and centralized control systems, ensuring the integrity of real-time operational data. The Transport Layer Security (TLS) 1.3 protocol, introduced in 2018, has been adopted for secure communication in power grids, safeguarding against data interception and tampering [restuccia2020low].
@@ 4.6 AI-Driven Incident Response
AI-driven incident response systems in OT environments leverage a combination of machine learning, anomaly detection, and automated decision-making algorithms to enhance cybersecurity resilience [nutalapati2024automated]. These systems employ advanced techniques like behavioral anomaly detection, where machine learning models continuously monitor device activity and network traffic for deviations from baseline patterns[he2020intelligent]. In addition to isolating compromised devices, they integrate with network segmentation protocols to minimize the spread of attacks. Sophisticated orchestration frameworks enable the automatic deployment of mitigation actions, such as blocking specific IP addresses or rerouting traffic, and rollback mechanisms use secure snapshots to restore devices to their pre-incident states [zhou2020unified]. In some cases, systems also use reinforcement learning to adapt and improve response strategies over time, making them more effective in handling novel, previously unseen attack vectors.
@@ 4.8 Integration of Cyber-Physical Security
Integrating cyber and physical security measures is essential for holistic OT protection [kure2018integrated]. Advanced systems now combine physical surveillance data, such as access logs, with cybersecurity monitoring to detect coordinated attacks.
@@
5 Performance Comparison: AI vs. Traditional OT Security Methods
The security of Operational Technology (OT) networks has traditionally relied on conventional techniques such as rule-based Intrusion Detection Systems (IDSs), signature-based detection, and network firewalls. These methods, while effective in identifying known threats, often struggle against evolving attack patterns and zero-day vulnerabilities. The increasing complexity of cyber threats targeting OT environments has led to the integration of artificial intelligence (AI)-driven security mechanisms. This section presents a quantitative comparison of traditional security approaches with AI-based techniques, demonstrating their relative effectiveness in OT networks.
Several studies have evaluated the performance of various cybersecurity strategies in OT environments, focusing on detection accuracy, false positive rates, and computational efficiency. Table 3 consolidates findings from multiple sources, including empirical evaluations using datasets such as KDD’99, NSL-KDD, and UNSW-NB15 [garcia2009anomaly, moustafa2016evaluation], as well as experimental validation on SCADA systems  and AI-driven security models for industrial control networks [vinayakumar2019deep, DBLP:journals/corr/abs-1807-07282].
Method Detection Accuracy (%) False Positives (%) Detection Time (ms) Adaptability
Rule-based IDS (Snort, Suricata) 80–85 10–15 150–200 Low adaptability, struggles with zero-day threats [garcia2009anomaly, moustafa2016evaluation]
Signature-based IDS (YARA, OSSEC) 86–90 8–12 120–180 Medium adaptability, detects known threats but limited against unknown attacks [garcia2009, inoue2017]
AI-based Anomaly Detection (ML/DL) 92–96 5–8 80–120 High adaptability, learns from new patterns and detects evolving threats [inoue2017, vinayakumar2019, DBLP:journals/corr/abs-1807-07282]
Hybrid AI + Signature-based IDS 95–98 3–5 60–100 Very High adaptability, combines AI-based learning with traditional detection methods [DBLP:journals/corr/abs-1807-07282, mitre2021]
Table 3: Performance Metrics of AI vs. Traditional OT Security Methods
The results indicate that AI-based models generally outperform traditional IDS techniques, achieving detection accuracies of up to 96%, with false positive rates reduced to below 8%. In comparison, rule-based IDS systems, despite their widespread use, exhibit lower detection accuracy and higher false positive rates, often exceeding 10%. The primary limitation of rule-based approaches is their reliance on predefined attack signatures, making them ineffective against zero-day attacks and adaptive threats [garcia2009anomaly]. Signature-based IDS solutions improve upon this by employing more extensive databases of known attack patterns, yet they still fall short in identifying novel attack behaviors, particularly in dynamic OT environments [inoue2017].
AI-driven intrusion detection, particularly deep learning-based models, addresses many of these limitations by identifying anomalous behaviors without requiring explicit attack signatures. These models analyze network traffic and system events in real-time, enabling faster detection while minimizing false positives. The results in Table 3 suggest that AI-based anomaly detection techniques achieve a balance between high detection accuracy and reduced computational overhead, with detection times averaging between 80–120 milliseconds, significantly lower than traditional methods [vinayakumar2019deep].
The adaptability of AI-driven security models is particularly relevant in OT networks, where attack strategies evolve rapidly. Unlike static rule-based IDS, machine learning-based approaches continuously refine their detection capabilities by learning from new attack patterns. However, AI models also introduce computational challenges, as deep learning-based security mechanisms require periodic retraining to maintain effectiveness. Furthermore, hybrid models that integrate AI-driven anomaly detection with traditional signature-based approaches have demonstrated superior performance, achieving detection rates of up to 98% while maintaining false positives as low as 3–5% [DBLP:journals/corr/abs-1807-07282]. This suggests that while standalone AI-based solutions provide significant improvements, a combined approach leveraging both AI and traditional techniques may offer the most effective defense against evolving cyber threats in OT environments.